OpenStudy (anonymous):

What is the equation of the line that passes through the point (4, 1) and has a slope of 5?

OpenStudy (anonymous):

@cammyabbo can you help me with this ?

OpenStudy (anonymous):

hi, yes i think so

OpenStudy (anonymous):

so we know we can set our midpoint equation equal to the (4,1)

OpenStudy (skullpatrol):

Try to help yourself first by telling us what you would try first.

OpenStudy (anonymous):

y = 5x - 3 y = 5x - 1 y = 5x + 3 y = 5x - 19 those are the options

OpenStudy (anonymous):

Y=mx+b......m is our slope and b is our y intercept

OpenStudy (anonymous):

we know our slope is 5

OpenStudy (anonymous):

so, so far the equation is y=5x

OpenStudy (skullpatrol):

Not, quite. So far the equation is $$y=5x +b$$

OpenStudy (anonymous):

right

OpenStudy (anonymous):

ok so whats the next step ?

OpenStudy (anonymous):

you know the line passes through (4,1)

OpenStudy (anonymous):

so 1=5(4)+b

OpenStudy (anonymous):

1=20+b

OpenStudy (anonymous):

your y-int is -19

OpenStudy (anonymous):

thats one of the options :)

OpenStudy (anonymous):

the equation of your line is y=5x-19

OpenStudy (anonymous):

so its the last one?

OpenStudy (anonymous):

all you did was plug (4,1) in the standard y=mx+b equation

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you sooooo much !!

OpenStudy (anonymous):

my pleasure

OpenStudy (anonymous):

Another way to do it is to use y-y1=m(x-x1) with m being your slope. This is just point slope form, so it's the same thing. You'd get y-1=5(x-4) so you'd have y-1=5x-20 and then y=5x-19. Not saying either way is better, but just in case you want this form too, that's it. :)